# HOG-Diff: Higher-Order Guided Diffusion for Graph Generation

### v1

@@ Abstract
Graph generation is a critical yet challenging task as empirical analyses require a deep understanding of complex, non-Euclidean structures. Although diffusion models have recently made significant achievements in graph generation, these models typically adapt from the frameworks designed for image generation, making them ill-suited for capturing the topological properties of graphs. In this work, we propose a novel Higher-order Guided Diffusion (HOG-Diff) model that follows a coarse-to-fine generation curriculum and is guided by higher-order information, enabling the progressive generation of plausible graphs with inherent topological structures. We further prove that our model exhibits a stronger theoretical guarantee than classical diffusion frameworks. Extensive experiments on both molecular and generic graph generation tasks demonstrate that our method consistently outperforms or remains competitive with state-of-the-art baselines. Our code is available at https://github.com/Yiminghh/HOG-Diff.
Graph Generation, Higher-order Networks, Score-based Generative Model, Topological Deep Learning
@@ 1 Introduction
Figure 1: Overivew of HOG-Diff. The dashed line above illustrates the classical generation process, where graphs quickly degrade into random structures with uniformly distributed entries. In contrast, as shown in the coloured region below, HOG-Diff adopts a coarse-to-fine generation curriculum based on the diffusion bridge, explicitly learning higher-order structures during intermediate steps with theoretically guaranteed performance.
Graphs provide an elegant abstraction for representing complex empirical phenomena by encoding entities as vertices and their relationships as edges, thereby transforming unstructured data into analyzable representations. Modelling the underlying distribution of graph-structured data is a crucial yet challenging task with broad applications, including social network analysis, motion synthesis, drug discovery, protein design, and urban planning (Zhu et al., 2022). The study of graph generation seeks to synthesize graphs that align with the observed distribution and traces back to seminal models of random network models (Erdős et al., 1960; Barabási & Albert, 1999). While these models offer foundational insights, they are often too simplistic to capture the complexity of graph distributions we encounter in practice.
@@
Recently, advances in generative models have leveraged the power of deep neural networks to significantly improve the ability to learn graph distributions. Notable approaches include models based on recurrent neural networks (RNNs) (You et al., 2018), variational autoencoders (VAEs) (Jin et al., 2018), and generative adversarial networks (GANs) (De Cao & Kipf, 2018; Martinkus et al., 2022). However, the end-to-end structure of these methods makes them hard to train. More recently, diffusion-based models have achieved remarkable success in image generation by learning a model to denoise a noisy sample (Ho et al., 2020; Song et al., 2021). With the advent of diffusion models, their applications on graphs with complex topological structural properties have recently aroused significant scientific interest (Niu et al., 2020; Jo et al., 2022; Vignac et al., 2023).
Despite these advances, existing graph generative models typically inherit the frameworks designed for image generation (Song et al., 2021), which fundamentally limits their ability to capture the intrinsic topological properties of networks. Notably, networks exhibit higher-order structures, such as motifs, simplices, and cells, which capture multi-way interactions and critical topological dependencies beyond pairwise relationships (Battiston et al., 2020; Zeng et al., 2024b; Papamarkou et al., 2024). These structures are vital for representing complex phenomena in domains like molecular graphs, social networks, and protein interactions. However, current methods are ineffective at modelling the topological properties of higher-order systems since learning to denoise the noisy samples does not explicitly preserve the intricate structural dependencies required for generating realistic graphs.
@@
Moreover, the image corrupted by Gaussian noise retains recognizable numerical patterns during the early and middle stages of forward diffusion. By contrast, the graph adjacency matrix quickly degrades into a dense matrix with uniformly distributed entries within a few diffusion steps. In addition, directly applying diffusion-based generative models to graph topology generation by injecting isotropic Gaussian noise to adjacency matrices is harmful as it destroys critical graph properties such as sparsity and connectivity. Lastly, such a framework should ensure equivariance111invariance as a particular special case, maintaining the learned distribution despite node index permutations, which is essential for robustness and capturing intrinsic graph distribution. Therefore, a graph-friendly diffusion process should also retain meaningful intermediate states and trajectories, avoid inappropriate noise addition, and ensure equivariance.
@@
Motivated by these principles and advances in topological deep learning (Hajij et al., 2022; Papamarkou et al., 2024), we propose the Higher-order Guided Diffusion (HOG-Diff) framework, illustrated in Fig. 1, to address the gaps in graph generation. HOG-Diff introduces a coarse-to-fine generation curriculum that enhances the model’s ability to capture complex graph properties by preserving higher-order topologies throughout the diffusion process. Specifically, we decompose the graph generation task into manageable sub-tasks, beginning by generating higher-order graph skeletons that capture core structures, which are then refined to include pairwise interactions and finer details, resulting in complete graphs with both topological and semantic fidelity. Additionally, HOG-Diff integrates diffusion bridge and spectral diffusion to ensure effective generation and adherence to the aforementioned graph generation principles. Our theoretical analysis reveals that HOG-Diff converges more rapidly in score matching and achieves sharper reconstruction error bounds than classical approaches, offering strong theoretical support for the proposed framework. Furthermore, our framework promises to enhance interpretability by enabling the analysis of different topological guides’ performance in the generation process. The contributions of this paper are threefold:
@@
  • Algorithmic: we introduce a coarse-to-fine graph generation curriculum guided by higher-order topological information using the OU diffusion bridge.
@@
  • Theoretical: our analysis reveals that HOG-Diff achieves faster convergence during score-matching and a sharper reconstruction error bound compared to classical methods.
@@
  • Experimental: extensive evaluations show that HOG-Diff achieves state-of-the-art graph generation performance across various datasets, highlighting the functional importance of topological guidance.
@@ 2 Preliminaries
Graphs are elegant and useful abstractions for various empirical objects; typically, they can be represented as $\bm{G}\triangleq(\bm{V},\bm{E},\bm{X})$. Here, $\bm{V}$ denote the node set, $\bm{E}\subseteq\bm{V}\times\bm{V}$, and $\bm{X}$ denotes the nodes feature matrix. However, many empirical systems exhibit group interactions that extend beyond simple pairwise relationships (Battiston et al., 2020). To capture these complex interactions, higher-order networks—such as hypergraphs, simplicial complexes, and cell complexes—offer more expressive alternatives by capturing higher-order interactions among entities (Papamarkou et al., 2024). Among these, cell complexes are fundamental in algebraic topology, offering a flexible generalization of pairwise graphs (Hatcher, 2001).
Definition 2.1 (Regular cell complex).
@@
    For each cell $x_{\alpha}$, the boundary $\partial x_{\alpha}$ is a union of finitely many cells, each having dimension less than that of $x_{\alpha}$.
@@
    Each cell $x_{\alpha}$ is homeomorphic to $\mathbb{R}^{n_{\alpha}}$, where $n_{\alpha}$ is dimension of $x_{\sigma}$.
@@
    (Regularity) For every $\alpha\in P_{\mathcal{S}}$, there exists a homeomorphism $\phi$ of a closed ball $\mathbb{B}^{n_{\alpha}}\subset\mathbb{R}^{n_{\alpha}}$ to the closure $\overline{x_{\sigma}}$ such that the restriction of $\phi$ to the interior of the ball is a homeomorphism onto $x_{\alpha}$.
From this definition, we can derive that $\mathcal{S}$ is the union of the interiors of all cells, i.e., $\mathcal{S}=\cup_{\alpha\in P_{\mathcal{S}}}int(x_{\alpha})$, where $int(x_{\alpha})$ denotes the interior of the cell $x_{\alpha}$. Intuitively, a cell complex can be constructed hierarchically through a gluing procedure. It begins with a set of vertices (0-cells), to which edges (1-cells) are attached by gluing the endpoints of closed line segments, thereby forming a graph. This process can be extended by taking a two-dimensional closed disk and gluing its boundary (i.e., a circle) to a simple cycle in the graph. While we typically focus on dimensions up to two, this framework can be further generalized by gluing the boundary of $n$-dimensional balls to specific $(n-1)$-cells in the complex. In this work, we also utilize simplicial complexes (SCs), a simpler and more restrictive subclass where 2-cells are limited to triangle shapes. A further introduction to higher-order networks can be found in App. B.
Figure 2: Visualization of Cell Complex Transformations. (a) An example graph. (b) Cell complex representation with corresponding homeomorphisms to closed balls for three example cells. (c) Black elements represent high-order structures extracted from the original graph through 2-cell filtering, while those in grey denote corresponding peripheral structures.
@@
A fundamental goal of generative models is to produce plausible samples from an unknown target data distribution $p(\mathbf{x}_{0})$. Score-based diffusion models (Song & Ermon, 2019; Song et al., 2021) achieve this by progressively corrupting the authentic data with noise and subsequently training a neural network to reverse this corruption process, thereby generating meaningful data from a tractable prior distribution, i.e., $\mathbf{x}_{generated}\sim p(\mathbf{x}_{0})$.
@@
where $\mathbf{f}_{t}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a vector-valued drift function, $g_{t}:[0,T]\to\mathbb{R}$ is a scalar diffusion coefficient, and $\mathbf{w}_{t}$ represents a Wiener process. Typically, $p(\mathbf{x}_{0})$ evolves over time $t$ from 00 to a sufficiently large $T$ into $p(\mathbf{x}_{T})$ through the SDE, such that $p(\mathbf{x}_{T})$ will approximate a tractable prior distribution, for example, a standard Gaussian distribution.
Starting from time $T$, $p(\mathbf{x}_{T})$ can be progressively transformed back to $p(\mathbf{x}_{0})$ by following the trajectory of the following reverse SDE (Anderson, 1982):
$\mathrm{d}\mathbf{x}_{t}=[\mathbf{f}_{t}(\mathbf{x}_{t})-g_{t}^{2}\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})}_{\text{score function}}]\mathop{}\!\mathrm{d}{t}+g_{t}\mathrm{d}\bar{\mathbf{w}},$ (2)
where $p_{t}(\cdot)$ denote the probability density function of $\mathbf{x}_{t}$ and $\bar{\mathbf{w}}$ is a reverse-time Wiener process. The score function is typically parameterized as a neural network $\bm{s}_{\bm{\theta}}(\mathbf{x}_{t},t)$ and trained using the score-matching technique as the loss function (Vincent, 2011):
$\begin{split}\ell(&\bm{\theta})\triangleq\mathbb{E}_{t,\mathbf{x}_{t}}\left[\omega(t)\left\|\bm{s}_{\bm{\theta}}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})\right\|^{2}\right]\\ &\propto\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{t}}\left[\omega(t)\left\|\bm{s}_{\bm{\theta}}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})\right\|^{2}\right],\end{split}$ (3)
where $\omega(t)$ is a weighting function, and the second expression is more commonly used as the conditional probability $p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ is generally accessible. Ultimately, the generation process is complete by first sampling $\mathbf{x}_{T}$ from a tractable prior distribution $p(\mathbf{x}_{T})\approx p_{prior}(\mathbf{x})$ and then generating $\mathbf{x}_{0}$ by numerically solving Eq. 2.
@@
Doob’s $h$-transform is a mathematical framework widely used to modify stochastic processes, enabling the process to satisfy specific terminal conditions. By introducing an $h$-function into the drift term of an SDE, this technique ensures that the process transitions to a predefined endpoint while preserving the underlying probabilistic structure. Specifically, given the SDE in Eq. 1, Doob’s $h$-transform alters the SDE to include an additional drift term, ensuring that the process reaches a fixed terminal point $t=T$. The modified SDE is expressed as:
@@
where $\bm{h}(\mathbf{x}_{t},t,\mathbf{x}_{T},T)=\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{T}|\mathbf{x}_{t})$. Crucially, the construction guarantees the conditional density satisfies $p(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})=1$ at $t=T$, ensuring that SDE terminates at the specified endpoint $\mathbf{x}_{T}$.

### v2

@@ Abstract
Graph generation is a critical yet challenging task as empirical analyses require a deep understanding of complex, non-Euclidean structures. Diffusion models have recently made significant achievements in graph generation, but these models are typically adapted from image generation frameworks and overlook inherent higher-order topology, leaving them ill-suited for capturing the topological properties of graphs. In this work, we propose Higher-order Guided Diffusion (HOG-Diff), a principled framework that progressively generates plausible graphs with inherent topological structures. HOG-Diff follows a coarse-to-fine generation curriculum guided by higher-order topology and implemented via diffusion bridges. We further prove that our model exhibits a stronger theoretical guarantee than classical diffusion frameworks. Extensive experiments on both molecular and generic graph generation tasks demonstrate that our method consistently outperforms or remains competitive with state-of-the-art baselines. Our code is available at https://github.com/Yiminghh/HOG-Diff.
@@ 1 Introduction
Figure 1: Overview of HOG-Diff. The dashed line above illustrates the classical generation process, where graphs quickly degrade into random structures with uniformly distributed entries. In contrast, as shown in the coloured region below, HOG-Diff adopts a coarse-to-fine generation curriculum based on the diffusion bridge, explicitly learning higher-order structures during intermediate steps with theoretical guarantees.
Graphs provide an elegant abstraction for representing complex systems by encoding entities as vertices and their relationships as pairwise edges. Yet mounting evidence shows that many empirical systems are governed by higher-order group interactions (Battiston et al., 2020). For instance, many molecules must assemble into specific structures to function properly, academic collaborations are most productive at specific group scales, and neural activity emerges from coordinated population dynamics (Papamarkou et al., 2024). Recent studies further reveal that higher-order interactions are ubiquitous and essential for understanding complex phenomena in domains as diverse as online social dynamics, ecological stability, and neuronal systems (Gardner et al., 2022).
Graph generation seeks to synthesize novel graphs faithful to the observed data distribution, with applications spanning drug discovery, motion synthesis, social network analysis, and urban planning (Zhu et al., 2022). However, modeling the underlying distribution of relational data remains challenging, partly because conventional approaches often overlook or implicitly learn data topology. Since higher-order structures are intrinsic to many empirical systems, explicitly incorporating them into generative frameworks promises to simplify generation and enhance realism. For instance, statistical analyses reveal that approved drug molecules occupy only a few hundred distinct higher-order structures (Ertl et al., 2025), in sharp contrast to the astronomically large chemical space of drug-like molecules ($10^{23}$$10^{60}$(Segler et al., 2018), suggesting that leveraging such topological guides could yield more efficient molecular design. Recent advances in topological deep learning (TDL) further reinforce this view, demonstrating that higher-order topology can substantially improve graph and molecular representation learning (Hajij et al., 2022; Papamarkou et al., 2024; Liu et al., 2024; Battiloro et al., 2025). Yet, despite its promise, higher-order topology has not been explicitly integrated into generative models, leaving a significant gap.
@@
Early work on graph generative models dates back to random network models (Barabási & Albert, 1999), which offer foundational insights but are too simplistic for capturing real-world graph distributions. Recent advances in generative models have leveraged the power of deep neural networks, significantly improving the ability to learn graph distributions. Notable approaches include models based on recurrent neural networks (RNNs) (You et al., 2018), variational autoencoders (VAEs) (Jin et al., 2018), and generative adversarial networks (GANs) (Martinkus et al., 2022). However, the end-to-end structure of these methods makes them hard to train. More recently, diffusion-based models have achieved remarkable success in image generation by learning a model to denoise a noisy sample (Ho et al., 2020; Song et al., 2021; Yue et al., 2024). With the advent of diffusion models, their applications on graphs with complex topological structural properties have recently aroused significant scientific interest (Niu et al., 2020; Jo et al., 2022; Vignac et al., 2023; Qin et al., 2025). Despite these advances, diffusion-based generative approaches are ineffective at modeling the topological properties of higher-order systems since learning to denoise the noisy samples does not explicitly preserve the intricate structural dependencies required for generating realistic graphs.
@@
Moreover, existing diffusion-based models for graphs typically inherit the frameworks designed for image generation (Song et al., 2021), which fundamentally limits their ability to capture the intrinsic topological properties of networks. Specifically, the image corrupted by Gaussian noise retains recognizable patterns during the early and middle stages of forward diffusion. By contrast, the graph adjacency matrix quickly degrades into a dense matrix with uniformly distributed entries within a few diffusion steps. In addition, applying diffusion-based generative models directly to graph topology generation by injecting isotropic Gaussian noise into adjacency matrices is detrimental, as it destroys essential graph properties such as sparsity and connectivity (Yang et al., 2023). Lastly, such a framework should ensure equivariance, with invariance as a special case, so that the learned distribution remains unchanged under node reordering, which is essential for robustness and capturing intrinsic graph distribution. Taken together, these challenges highlight the need for a graph-friendly diffusion framework that explicitly learns higher-order topology, preserves meaningful intermediate states and trajectories, avoids inappropriate noise addition, and ensures equivariance.
@@
Motivated by these principles, we propose the Higher-order Guided Diffusion (HOG-Diff) framework (see Fig. 1), a coarse-to-fine generation curriculum that preserves higher-order topologies throughout the diffusion process to better capture complex graph structures. Specifically, we decompose the graph generation task into manageable sub-tasks, beginning by generating higher-order skeletons that capture core structures, which are then refined to include pairwise interactions and finer details, ultimately synthesizing complete graphs with both topological and semantic fidelity. Additionally, HOG-Diff integrates the generalized Ornstein-Uhlenbeck (OU) bridge in the spectral domain to ensure effectiveness and adherence to the aforementioned graph generation principles. Our theoretical analysis reveals that HOG-Diff converges more rapidly in score matching and achieves sharper reconstruction error bounds than classical approaches, offering strong theoretical support for the proposed framework. Furthermore, our framework promises to enhance interpretability by enabling the analysis of different topological guides’ performance in the generation process. The contributions of this paper are fourfold:
  • We introduce cell complex filtering to extract higher-order skeletons as valuable generation guides.
@@
  • We propose a principled coarse-to-fine graph generation framework guided by higher-order topological information and implemented via the generalized OU diffusion bridge.
@@
  • We theoretically show that HOG-Diff achieves faster convergence during score-matching and a sharper reconstruction error bound compared to classical diffusion models.
@@
  • Extensive evaluations on molecules and generic graph datasets show that HOG-Diff achieves state-of-the-art performance, highlighting the practical importance of topological guidance.
@@ 2 Preliminaries
Graphs are elegant and useful abstractions for various empirical objects. Formally, a graph can be represented as $\bm{G}\triangleq(\bm{V},\bm{E},\bm{X})$, where $\bm{V}$ denotes the node set, $\bm{E}\subseteq\bm{V}\times\bm{V}$ the edges, and $\bm{X}$ the nodes feature matrix. However, many empirical systems exhibit group interactions that go beyond simple pairwise relationships (Battiston et al., 2020). To capture these complex interactions, higher-order networks—such as hypergraphs, simplicial complexes (SCs), and cell complexes (CCs)—offer more expressive alternatives by capturing higher-order interactions among multiple entities (Papamarkou et al., 2024). Among these, cell complexes are fundamental in algebraic topology, offering a flexible generalization of pairwise graphs (Hatcher, 2001).
Definition 1 (Regular cell complex).
@@
    Each cell $x_{\alpha}$ is homeomorphic to $\mathbb{R}^{n_{\alpha}}$, where $n_{\alpha}=\dim(x_{\alpha})$ denotes the dimension of $x_{\alpha}$.
@@
    For each cell $x_{\alpha}$, the boundary $\partial x_{\alpha}$ is a finite union of cells of dimension less than $\dim(x_{\alpha})$.
@@
    For every $\alpha\in P_{\mathcal{S}}$, there exists a homeomorphism $\phi_{\alpha}$ of a closed ball $\mathbb{B}^{n_{\alpha}}\subset\mathbb{R}^{n_{\alpha}}$ to the closure $\overline{x_{\alpha}}$ such that the restriction of $\phi_{\alpha}$ to the interior of the ball is a homeomorphism onto $x_{\alpha}$.
Lifting: From Graphs to Cell Complexes
A cell complex can be constructed hierarchically through a gluing procedure, which is known as lifting. It begins with a set of vertices (0-cells), to which edges (1-cells) are attached by gluing the endpoints of closed line segments, thereby forming a graph. This process can be extended by taking a two-dimensional closed disk and gluing its boundary to a simple cycle in the graph, see Fig. 2 for illustration. While we typically focus on dimensions up to two, this framework can be further generalized by gluing the boundary of $n$-dimensional balls to specific (n1)(n-1)-cells in the complex.
From the definition, we can derive that the cell complex $\mathcal{S}$ is the union of the interiors of all cells. In this work, we also consider simplicial complexes (SCs), a class of topological spaces represented by finite sets of elements that are closed under the inclusion of subsets. Intuitively, SCs can be viewed as a more constrained subclass of cell complexes, where 2-cells are limited to triangle shapes. A comprehensive introduction to higher-order networks can be found in Sec. B.1.
@@
A fundamental goal of generative models is to produce plausible samples from an unknown target data distribution $p(\mathbf{x}_{0})$. Score-based diffusion models (Song & Ermon, 2019; Song et al., 2021) achieve this by progressively corrupting the authentic data with noise and subsequently training a neural network to reverse this corruption process, thereby generating meaningful data from a tractable prior distribution, i.e., $\mathbf{x}_{\mathrm{generated}}\sim p(\mathbf{x}_{0})$.
@@
where $\mathbf{f}_{t}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a vector-valued drift function, $g_{t}:[0,T]\to\mathbb{R}$ is a scalar diffusion coefficient, and $\mathbf{w}_{t}$ represents a Wiener process. Typically, $p(\mathbf{x}_{0})$ evolves over time $t$ from 0 to a sufficiently large $T$ into $p(\mathbf{x}_{T})$ through the SDE, such that $p(\mathbf{x}_{T})$ will approximate a tractable prior distribution, for example, a standard Gaussian distribution.
Starting from time $T$, $p(\mathbf{x}_{T})$ can be progressively transformed back to $p(\mathbf{x}_{0})$ by following the trajectory of the reverse-time SDE $\mathrm{d}\mathbf{x}_{t}=[\mathbf{f}_{t}(\mathbf{x}_{t})-g_{t}^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})]\mathop{}\!\mathrm{d}{\bar{t}}+g_{t}\mathrm{d}\bar{\mathbf{w}}_{t}$ (Anderson, 1982), where $p_{t}(\cdot)$ denotes the probability density function of $\mathbf{x}_{t}$ and $\bar{\mathbf{w}}$ is a reverse-time Wiener process.
The term $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$, known as the score function, is typically parameterized by a neural network $\bm{s}_{\bm{\theta}}(\mathbf{x}_{t},t)$ and trained using the conditional score-matching loss function (Vincent, 2011):
$\ell(\bm{\theta})\triangleq\mathbb{E}_{t,\mathbf{x}_{t}}\left[\omega(t)\left\|\bm{s}_{\bm{\theta}}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})\right\|^{2}\right]\propto\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{t}}\left[\omega(t)\left\|\bm{s}_{\bm{\theta}}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})\right\|^{2}\right],$
where $\omega(t)$ is a weighting function. The second expression is more commonly used since the conditional probability $p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ is generally accessible. Ultimately, the generation process is complete by first sampling $\mathbf{x}_{T}$ from a tractable prior distribution $p(\mathbf{x}_{T})\approx p_{\mathrm{prior}}(\mathbf{x})$ and then generating $\mathbf{x}_{0}$ by numerically solving the reverse-time SDE.
@@
Doob’s $h$-transform is a mathematical framework widely used to modify stochastic processes, enabling the process to satisfy specific terminal conditions. By introducing a $h$-function into the drift term of an SDE, this technique ensures that the process transitions to a predefined endpoint while preserving the underlying probabilistic structure. Specifically, given the SDE in Eq. 1, Doob’s $h$-transform alters the SDE to include an additional drift term, ensuring that the process reaches a fixed terminal state at $t=T$. The modified SDE is expressed as:
@@
where $\bm{h}(\mathbf{x}_{t},t,\mathbf{x}_{T},T)=\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{T}|\mathbf{x}_{t})$. Crucially, the construction drives the diffusion process towards a Dirac distribution at $\mathbf{x}_{T}$, i.e., $\lim_{t\to T}p(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})=\delta(\mathbf{x}_{t}-\mathbf{x}_{T})$.